### v2

# Identification of fixations and saccades in eye-tracking data using adaptive threshold-based methods

Abstract
Properties of ocular fixations and saccades are highly stochastic during many experimental tasks, and their statistics are often used as proxies for various aspects of cognition. Although distinguishing saccades from fixations is not trivial, experimentalists generally use common ad-hoc thresholds in detection algorithms. This neglects inter-task and inter-individual variability in oculomotor dynamics, and potentially biases the resulting statistics. In this article, we introduce and evaluate an adaptive method based on a Markovian approximation of eye-gaze dynamics, using saccades and fixations as states such that the optimal threshold minimizes state transitions. Applying this to three common threshold-based algorithms (velocity, angular velocity, and dispersion), we evaluate the overall accuracy against a multi-threshold benchmark as well as robustness to noise. We find that a velocity threshold achieves the highest baseline accuracy (90-93%) across both free-viewing and visual search tasks. However, velocity-based methods degrade rapidly under noise when thresholds remain fixed, with accuracy falling below 20% at high noise levels. Adaptive threshold optimization via $K$-ratio minimization substantially improves performance under noisy conditions for all algorithms. Adaptive dispersion thresholds demonstrate superior noise robustness, maintaining accuracy above 81% even at extreme noise levels ($\sigma=50$ px), though a precision-recall trade-off emerges that favors fixation detection at the expense of saccade identification. In addition to demonstrating our parsimonious adaptive thresholding method, these findings provide practical guidance for selecting and tuning classification algorithms based on data quality and analytical priorities.
keywords:
Eye tracking, Fixation–saccades categorization, Adaptive thresholding, Noise robustness, Visual search
1 Introduction
The human visual system orchestrates a sophisticated interplay of eye movements, fixations, saccades, smooth pursuits, and vergence—that reveals intricate patterns of attention, cognition, and neurological function [Duchowski2017]. Eye-tracking technology, utilizing corneal reflections and pupil dynamics at sampling rates often exceeding 1,000 Hz, captures these movements with high precision, mapping foveal gaze to elucidate visual perception [68d3bfb814cc4a61aae0131dd135c9ec]. Fixations, stable gazes lasting 200–400 ms to process detailed visual information, and saccades, rapid eye movements reaching velocities up to 500/s, form the cornerstone of applications across diverse fields, including psychology, neuroscience, clinical diagnostics, education, marketing, and human-computer interaction [10.1145/355017.355028]. From early diagnosis of neurological disorders such as Alzheimer’s and schizophrenia [korda2018automatic], to enhancing educational tools through gaze-based learning analytics [busjahn2014eye], to optimizing consumer behavior analysis in visual marketing [MKT-011], eye-tracking has transformative potential. Yet, accurately distinguishing fixations from saccades remains a central challenge, as traditional algorithms struggle with noise, blinks, and behavioral variability across diverse visual stimuli [schweitzer2020adaptive].
Conventional algorithms, such as Velocity-Threshold Identification (I-VT), Angular Velocity-Threshold Identification (I-AVT), and Dispersion-Threshold Identification (I-DT), rely on fixed velocity (e.g., 20–40°/s for I-VT) or spatial dispersion thresholds to classify eye movements [10.1145/355017.355028]. While computationally efficient and suitable for real-time applications, these methods are sensitive to noise and often fail to detect auxiliary events like microsaccades or post-saccadic oscillations (glissades), limiting their robustness in complex datasets [martinezs, nystrom2010adaptive]. Recent advancements have introduced more sophisticated approaches to address these limitations. For instance, Larsson et al. developed robust segmentation techniques for fixations and smooth pursuits in high-speed eye-tracking data, outperforming velocity- and dispersion-based methods [LARSSON2015145]. Zemblys et al. leveraged machine learning, specifically random forest classifiers, to automate event detection with accuracy comparable to manual coding [zemblys]. Additionally, methods like hidden Markov models (HMM) and minimum spanning trees offer nuanced classification by modeling temporal and spatial gaze patterns [10.1145/355017.355028]. Despite these innovations, traditional algorithms remain prevalent due to their simplicity and versatility, necessitating rigorous evaluation to optimize their performance across varied tasks [10.1145/355017.355028]. Recent comprehensive evaluations have demonstrated that algorithm choice significantly impacts classification outcomes [Andersson2017OneAlgorithm], motivating the need for systematic comparison studies.
This study evaluates the efficacy of I-VT, I-AVT, and I-DT for classifying fixations and saccades in eye-tracking data collected from free viewing and visual search tasks, and introduces an adaptive method for threshold selection based on minimization of state transitions, using a quantity we term the $K$-ratio. We then perform a Gaussian noise analysis to evaluate robustness (e.g., detecting performance degradation at noise levels $\geq$ 2 pixels).The analysis focuses on implementation difficulty, computational complexity, usability, and classification effectiveness, addressing the question: “Which algorithms are best suited for classifying eye-tracking features and movements?” Our findings aim to provide actionable insights for improving classification accuracy, advancing applications in cognitive research, clinical diagnostics, and assistive technologies.
2 Data and methods
2.1 “Pure” and “noisy” eye-gaze trajectories
[…]
Two paradigms were used for the same 15 participants. In a free-viewing condition (Random Pixel), participants viewed unstructured pixel fields (3 trials, 60 s each; $\sim$180,000 samples per participant, $\sim$2.7 M total in this subset). In a visual search condition (Waldo), participants located targets in complex scenes from Where’s Waldo? (9 trials, 45 s each; $\sim$405,000 samples per participant, $\sim$6 M total in this subset) [bookwaldo]. Each dataset comprised ten columns per sample at 1000 Hz: timestamp; left and right $x,y$ gaze in pixels; left and right pupil size; and event flags. See Tab. 1.
Table 1: Comparison of random pixel and Waldo task datasets.
Characteristic Random pixel task Waldo task
Number of participants 15 15
Data points per participant $\sim$180,000 $\sim$405,000
Total data points 2,700,252 $\sim$6,000,000
Row count range 180,003–180,044 405,088–405,163
Number of columns 10 10
Sampling rate 1000 Hz 1000 Hz
Approximate duration 3 minutes 6.75 minutes
Binocular coordination was assessed via Pearson correlation and linear regression on left/right eye X/Y coordinates, with fixation/saccade agreement calculated as percentage overlap. In the Random Pixel dataset, binocular coordination was high: Pearson correlations between left and right eyes were (r=0.9948) for the (x)-coordinate and (r=0.9748) for the (y)-coordinate. Framewise agreement between the two eyes’ event labels was 98.30% for fixations and 98.34% for saccades. Across the full datasets, the overall event distribution was dominated by fixations ($\sim 92\%)$. In the rest of this article, we use the left eye only for the X/Y coordinates.
The EyeLink software uses a combination of velocity and acceleration thresholds to classify saccades and fixations. We used the default settings (in degrees of visual angle: 30/s velocity, 8000/sec2 acceleration) to extract ground truth labels per frame.
To study the effect of noise in the classification algorithm, we augmented our data sample by generating a superposition of the collected trajectories with a tunable level of noise. These “noisy” trajectories were then analyzed additionally to the collected eye-gaze trajectories. We restricted to the case of Gaussian noise, with $\varepsilon_{x,i}$ and $\varepsilon_{y,i}$, added to each observation $i$, at the location $(x_{i},y_{i})$:
$P^{\text{(noisy)}}=\{(x_{i}^{\text{(noisy)}},y_{i}^{\text{(noisy)}})\}=\{(x_{i}+\varepsilon_{x,i},y_{i}+\varepsilon_{y,i})\}\,$, (1)
where $\varepsilon_{x,i},\varepsilon_{y,i}\sim\mathcal{N}(0,\sigma^{2})$ are normally distributed with zero mean and a tunable standard deviation $\sigma$.
2.2 Classification algorithms
We introduce adaptive versions of the three classification algorithms mentioned above, I-VT, I-AVT, and I-DT [10.1145/355017.355028]. The I-VT algorithm uses the instantaneous velocity $V_{i}$, which is the velocity calculated between successive measurements, namely:
$V_{i}=\frac{\sqrt{(x_{i+1}-x_{i})^{2}+(y_{i+1}-y_{i})^{2}}}{t_{i+1}-t_{i}},$ (2)
where $x_{i},y_{i}$ are coordinates at timestamp $t_{i}$.
The I-AVT method uses the effective angular velocity, which combines the scalar velocity and the directional persistence, namely
$V^{\text{eff}}_{i}=V_{i}\cdot\cos(\theta_{i}-\theta_{i-1}),$ (3)
where $\theta_{i}$ is the angle of the displacement vector with the positive horizontal axis:
$\theta_{i}=\tan^{-1}\left(\frac{y_{i+1}-y_{i}}{x_{i+1}-x_{i}}\right)\;.$ (4)
The I-DT algorithm identifies fixations based on the spatial dispersion of gaze points within a moving time window. It uses two parameters, a minimum time window span $T_{\mathrm{min}}$ and a dispersion threshold $D_{max}$. The algorithm starts at the beginning of the time series and searches for the first index $i$ where
$\exists k>i:\;(t_{k}-t_{i}>T_{\mathrm{min}})\land(D_{i}^{k}<D_{\mathrm{max}})\;,$ (5)
where $D_{i}^{k}$ is the dispersion within a time span $t_{i},t_{k}]$, combining the dimensions of the smallest rectangle that encompasses all recorded gaze locations between $t_{i}$ and $t_{k}$:
$D_{i}^{k}=\max_{j\in[i,k]}(x_{j})-\min_{j\in[i,k]}(x_{j})+\max_{j\in[i,k]}(y_{j})-\min_{j\in[i,k]}(y_{j})\;.$ (6)
All data points from index $i$ up to the largest index $k$ satisfying the condition in equation (5) are then labelled as a fixation. The next iteration of the algorithm then begins the search from index $k+1$.
2.3 Optimization through a statistical metric
In this section, we introduce a metric we call $K$-ratio, which compares the empirical number of fixation-saccade transitions observed in an eye-gaze trajectory with the number of transitions between saccades and fixations that would be expected if data points simply alternated randomly between being fixations and saccades:
$\text{$K$-ratio}=\frac{p_{\mathrm{empirical}_{F\to S}}}{p_{\mathrm{independentF\rightarrow S}}},$ (7)
Figure 1: Example of scanpaths with overlaid classifications (orange = fixation, blue = saccade). Top row: Random Pixel task; bottom row: Waldo task. Columns (left to right): raw gaze points, I-VT, I-AVT, and I-DT results.
If the $K$-ratio is lower than one, it means that the labels tend to agglomerate along the sample, meaning, in our case, that data points labelled as fixations (resp. saccades) tend to follow other data points labelled as fixations (resp. saccades).
If the probability to transition from one state to another is independent of the current state and independent of time, then the expected transition probability from a label $\ell_{\bar{V}_{i}}(V_{j})=F$ to $\ell_{\bar{V}_{i}}(V_{j+1})=S$ would be:
$p_{\mathrm{independent}_{F\to S}}=Pr\left(\ell_{\bar{V}_{i}}(V_{j+1})=S\right)Pr\left(\ell_{\bar{V}_{i}}(V_{j})=F\right)=n_{S}(1-n_{S}),$ (8)
where $n_{S}=N_{S}/N$, with $N_{S}$ the number of occurrences of the label $S$ and $N$ the total number of observations.
However, if the probability of transition from one state to another depends on the current state, this separation is not possible. The $K$-ratio is then
$\text{$K$-ratio}=\frac{Pr\left(\ell_{\bar{V}_{i}}(V_{j+1})=S,\ell_{\bar{V}_{i}}(V_{j})=F\right)}{Pr\left(\ell_{\bar{V}_{i}}(V_{j+1})=S\right)Pr\left(\ell_{\bar{V}_{i}}(V_{j})=F\right)}.$ (9)
Again, assuming stationarity, this can be written in terms of empirical quantities
$\text{$K$-ratio}=\frac{n_{F\rightarrow S}}{n_{S}(1-n_{S})}\;,$ (10)
where $n_{F\rightarrow S}$ is the number of fixation-to-saccade transitions divided by the number of data points.
By minimizing this ratio, thresholds used to label fixations and saccades are optimized towards keeping these two states separated in time.
While another kind of state-space modelling (i.e. hidden Markov model) has been used to adaptively identify fixations [10.1145/355017.355028], and velocity-adaptive approaches have been proposed [nystrom2010adaptive], these require detailed modeling of the velocity distributions of fixations and saccades. In comparison, the $K$-ratio is parsimonious and can be applied to any threshold-based algorithm.
Figure 2: $K$-ratio curves used for threshold optimization (top: Random Pixel; bottom: Waldo). Minima define the task-specific thresholds for I-VT, I-AVT, and I-DT. The respective $K$-ratio for each of the algorithms for both task are tabulated in a supplementary section at the end of the paper. See Appendix Tab. S2. The $T_{\mathrm{min}}$ value used for I-DT is 50ms.
3 Results
Figure 3: Performance metrics at zero noise (left: Random Pixel; right: Waldo). Bars display overall accuracy, class-wise precision/recall/F1, and predicted fixation proportion for each algorithm.
We evaluated I-VT, I-AVT, and I-DT on the Random Pixel (free viewing) and Waldo (visual search) datasets recorded at 1000 Hz, using the EyeLink event labels as frame-wise ground truth. The three classifiers were applied to the same recordings and compared in terms of threshold optimization, frame-wise accuracy, class-wise metrics, and robustness to added Gaussian noise. Frame-wise ground truth was provided by the standard EyeLink parser (30/s velocity; 8000/s2 acceleration).
Figure 1 shows representative scanpaths with overlaid classifications for both tasks. I-VT produces compact fixation clusters with clearly delineated saccade segments. I-AVT tends to generate long fixation runs with relatively few saccade segments, while I-DT yields the most fixation-dominated segmentation, often merging brief saccades into extended fixation episodes. These qualitative patterns reflect the underlying class distributions and confusion trends that appear in the aggregate metrics. Thresholds were selected by minimizing the $K$-ratio for each algorithm and task. The $K$-ratio curves for the Random Pixel and Waldo datasets are shown in Fig. 2, and the corresponding optimal thresholds and $K$-ratio values are summarized in Appendix Tab. S2. In both tasks, the $K$-ratio decreases sharply at low thresholds and then levels off into a broad minimum. The Waldo task exhibits wider minima than the Random Pixel task, consistent with more heterogeneous oculomotor dynamics during structured visual search.
Overall frame-wise accuracy and class-wise precision, recall, and F1-scores for fixations and saccades are summarized visually in Fig. 3. All algorithms slightly over-classify fixations compared to the ground truth proportion of 92%. Across both tasks, I-VT attains the highest overall accuracy and preserves near-complete recovery of fixations, while providing the strongest compromise between saccade precision and recall. I-AVT maintains high fixation performance but under-detects saccades, leading to low saccade F1-scores despite its use of directional information. I-DT shows a pronounced fixation bias, with the highest fixation proportions and low saccade recall, reflecting the tendency of the dispersion criterion to group short saccades into longer fixation segments. Taken together, the results indicate that I-VT offers the most balanced performance among the three methods when evaluated against the EyeLink labels.
Figure 4: Robustness analysis of fixation classification algorithms under Gaussian spatial noise ($\sigma=0$-5050 pixels). Top row: Evolution of adaptive thresholds for I-VT, I-AVT, and I-DT algorithms across noise levels for Random Pixel and Waldo tasks. Bottom row: Classification accuracy comparing fixed thresholds (optimized at $\sigma=0$) versus adaptive thresholds (re-optimized at each noise level). Solid lines represent original (fixed) thresholds; dashed lines represent adaptive thresholds.
The effect of added Gaussian noise ($\sigma$ = 0-50 pixels) was examined both with the original threshold (rapid degradation) – cf. Tabs. S3-S6 in the supplementary section – and with thresholds re-optimized at each noise level via the $K$-ratio. See Fig. 4.
Under clean recording conditions, I-VT yields the strongest agreement with the reference parser. When spatial noise is present, I-VT and I-AVT showed a progressive drift toward saccade-dominated classifications as noise increased, driven by inflated velocity and angular-velocity estimates. I-DT preserved more stable fixation proportions at intermediate noise levels but also deteriorated at higher amplitudes. With adaptive thresholding, I-DT shows the greatest stability in event proportions, while velocity-based algorithms remain markedly more sensitive to recording precision. However, in all cases, the $K$-ratio-based threshold at least partially mitigated against noise.
4 Discussion and conclusion
This study compared three threshold-based classifiers, I-VT, I-AVT, and I-DT, on free viewing (Random Pixel) and visual search (Waldo) eye-tracking data, using EyeLink labels as a reference. Across both tasks, I-VT provided the highest overall agreement with the benchmark and the most favorable balance between fixation and saccade classification. Task structure influenced the optimal thresholds obtained by $K$-ratio minimization. The Waldo dataset required higher velocity and angular-velocity thresholds than the Random Pixel dataset, which reflects more variable and goal-directed oculomotor dynamics in visual search. At the same time, both datasets were dominated by fixations, with saccades forming only a small fraction of samples. This class imbalance means that accuracy alone is not a sufficient measure of performance. The class-wise metrics in the Results show that a classifier can achieve relatively high accuracy while failing to identify many saccades. This supports the use of F1-scores, confusion matrices, and explicit event proportions, in line with earlier discussions of evaluation practice in eye movement classification [zemblys].
The $K$-ratio criterion provided a simple, task-specific method to tune thresholds without explicit modelling of the underlying velocity distributions. By minimizing the discrepancy between empirical and independent transition probabilities, it favored segmentation in which fixations and saccades formed temporally coherent states. This approach is complementary to more elaborate state-space models, such as hidden Markov models, which require stronger assumptions about observation statistics [10.1145/355017.355028]. Here, the $K$-ratio produced consistent thresholds across tasks and offered a transparent way to adapt classical I-VT, I-AVT, and I-DT schemes to raw eye-tracking data.
The noise analysis confirmed that all three algorithms are sensitive to measurement noise. Even modest Gaussian perturbations led to a marked reduction in agreement with the EyeLink labels and to substantial shifts in event composition. In all cases, re-optimizing thresholds by $K$-ratio mitigated, but did not entirely prevent, the loss of performance. These findings agree with previous work emphasizing the importance of careful preprocessing and denoising in eye-tracking analysis [niehorster2020noise, niehorster2017denoising, LARSSON2015145]. From an applications perspective, the results suggest that I-VT remains a reasonable default when high-quality data are available, and the goal is to obtain a reliable first-order segmentation of fixations and saccades. I-AVT, despite its directional component, offered limited benefit in this context because of its strong under-detection of saccades. I-DT may still be useful in settings where long, stable fixations are of primary interest, such as some reading and static-scene paradigms [68d3bfb814cc4a61aae0131dd135c9ec], but it is less suitable for dynamic visual search. In all cases, practitioners should be aware that fixed thresholds calibrated on one task or recording setup may not transfer to others, and that auxiliary events such as microsaccades, smooth pursuits, and glissades remain largely unmodeled [martinezs].
There are several natural directions for future work. One is to combine $K$-ratio tuning with more advanced denoising and interpolation schemes, which could extend the useful operating range of classical algorithms in noisy or mobile settings. Another is to embed $K$-ratio or related transition statistics into learning-based classifiers, for example, as regularization terms or diagnostic tools for model calibration [FRANCESCHIELLO2022106929, zemblys]. A third is to pursue more systematic study of inter-individual differences, with thresholds optimized at the individual level, to clarify how much of the observed variability is due to genuine oculomotor differences and how much is an artifact of algorithm choice [HESSELS2019100710].
In conclusion, I-VT, I-AVT, and I-DT remain useful tools for eye movement classification, but their performance depends strongly on task characteristics, threshold selection, and data quality. The $K$-ratio provides a simple and general way to adapt these methods to new datasets by minimizing unnecessary state transitions. Researchers should choose algorithms and thresholds in light of specific analytical goals, apply robust preprocessing, and be cautious in interpreting fine-grained differences in eye movement metrics when classification uncertainty is high. The development of standardized analysis protocols [5523936] remains an important goal for the field.
Declarations
  • Conflict of interest/Competing interests: The authors have no competing interests to declare that are relevant to the content of this article.
  • Data availability: Data sets used during the current study are available from the corresponding author on reasonable request.
Acknowledgements
The authors gratefully acknowledge support from the Research Council of Norway through the “VirtualEye” project (Ref. 335940-FORSKER22).
Supplementary results
In this supplementary section, we provide detailed results from the noise robustness analysis. Table S2 reports the task-specific thresholds obtained by $K$-ratio minimization under noise-free conditions. These values served as the fixed (non-adaptive) baseline thresholds in subsequent experiments.
Table S2: $K$-ratio-optimized thresholds for the three algorithms (zero added noise).
Algorithm Random Pixel Waldo
Threshold $K$-ratio Threshold $K$-ratio
I-VT 5.27 px/s 0.189 6.57 px/s 0.179
I-AVT 5.31 rad/s 0.190 6.75 rad/s 0.181
I-DT 10.20 px 0.020 4.98 px 0.029
Table S3: Classification degradation with fixed thresholds under increasing noise for I-VT algorithm. Note: Agr. = agreement with noise level 0 ($\sigma=0$) (%); Fix/Sac = sample counts; $\Delta$% = change in fixation proportion from baseline.
Random Pixel Waldo
$\sigma$ Agr. Fix Sac Fix% $\Delta$% Agr. Fix Sac Fix% $\Delta$%
0 100.0 80478 5110 94.0 0.0 100.0 47895 2104 95.8 0.0
1 98.5 79814 5774 93.3 0.8 99.3 47758 2241 95.5 0.3
2 80.4 64617 20971 75.5 18.5 91.4 43965 6034 87.9 7.9
5 27.9 19694 65894 23.0 71.0 36.8 16809 33190 33.6 62.2
10 12.0 5605 79983 6.6 87.5 13.7 5048 44951 10.1 85.7
30 6.7 708 84880 0.8 93.2 5.3 596 49403 1.2 94.6
40 6.4 383 85205 0.4 93.6 4.8 315 49684 0.6 95.2
50 6.2 245 85343 0.3 93.7 4.6 222 49777 0.4 95.4
Table S4: I-VT: Original vs. Adaptive threshold performance across noise levels. Note: Acc = Accuracy; F1F = F1-score (Fixation); F1S = F1-score (Saccade).
Random Pixel Waldo
Original Adaptive Original Adaptive
$\sigma$ Acc F1F F1S Acc F1F F1S Acc F1F F1S Acc F1F F1S
0 0.90 0.94 0.51 0.90 0.94 0.51 0.93 0.96 0.54 0.93 0.96 0.54
1 0.90 0.94 0.52 0.89 0.94 0.44 0.93 0.96 0.56 0.93 0.96 0.53
2 0.81 0.89 0.43 0.88 0.93 0.38 0.87 0.93 0.44 0.92 0.95 0.45
5 0.38 0.46 0.29 0.78 0.87 0.28 0.39 0.50 0.22 0.80 0.89 0.25
10 0.22 0.16 0.27 0.67 0.79 0.24 0.19 0.18 0.21 0.71 0.82 0.19
30 0.17 0.02 0.27 0.69 0.81 0.19 0.12 0.02 0.21 0.70 0.82 0.16
40 0.16 0.01 0.27 0.69 0.81 0.19 0.12 0.01 0.21 0.70 0.82 0.16
50 0.16 0.01 0.27 0.69 0.81 0.19 0.12 0.01 0.21 0.70 0.82 0.16
Table S5: I-AVT: Original vs. Adaptive threshold performance across noise levels. See Tab. S4.
Random Pixel Waldo
Original Adaptive Original Adaptive
$\sigma$ Acc F1F F1S Acc F1F F1S Acc F1F F1S Acc F1F F1S
0 0.84 0.94 0.46 0.89 0.94 0.46 0.89 0.96 0.53 0.90 0.96 0.53
1 0.89 0.94 0.45 0.89 0.94 0.43 0.93 0.96 0.52 0.93 0.96 0.53
2 0.87 0.92 0.40 0.87 0.93 0.36 0.90 0.94 0.43 0.91 0.95 0.42
5 0.60 0.72 0.27 0.79 0.88 0.21 0.59 0.72 0.21 0.81 0.89 0.20
10 0.39 0.49 0.26 0.68 0.80 0.20 0.36 0.47 0.20 0.76 0.86 0.14
30 0.24 0.21 0.27 0.42 0.52 0.26 0.20 0.20 0.20 0.41 0.53 0.19
40 0.22 0.16 0.27 0.36 0.44 0.26 0.18 0.15 0.20 0.34 0.43 0.20
50 0.21 0.13 0.27 0.31 0.35 0.27 0.17 0.12 0.20 0.29 0.37 0.20
Table S6: I-DT: Original vs. Adaptive threshold performance across noise levels. See Tab. S4.
Random Pixel Waldo
Original Adaptive Original Adaptive
$\sigma$ Acc F1F F1S Acc F1F F1S Acc F1F F1S Acc F1F F1S
0 0.90 0.93 0.42 0.87 0.93 0.44 0.94 0.95 0.48 0.91 0.95 0.54
1 0.87 0.93 0.42 0.87 0.93 0.37 0.91 0.95 0.47 0.91 0.95 0.49
2 0.85 0.92 0.40 0.86 0.92 0.31 0.89 0.94 0.43 0.90 0.95 0.36
5 0.63 0.75 0.30 0.85 0.92 0.11 0.63 0.76 0.24 0.90 0.94 0.20
10 0.34 0.40 0.27 0.85 0.92 0.11 0.32 0.41 0.21 0.89 0.94 0.14
30 0.18 0.07 0.27 0.84 0.91 0.08 0.14 0.07 0.21 0.88 0.94 0.09
40 0.17 0.04 0.27 0.82 0.90 0.10 0.13 0.04 0.21 0.86 0.93 0.10
50 0.17 0.03 0.27 0.81 0.90 0.10 0.12 0.02 0.21 0.85 0.92 0.10
Figure S5: F1 scores for fixation and saccade classification across noise levels. Top row: F1-score for fixation detection. Bottom row: F1-score for saccade detection. Solid lines indicate original thresholds; dashed lines indicate adaptive thresholds.
Figure S6: Precision and recall for fixation detection across noise levels. Top row: Precision. Bottom row: Recall. Adaptive thresholds maintain higher precision at the cost of reduced recall at extreme noise levels.
Table S3 shows how classification performance degrades when optimal noise-free thresholds are held fixed under increasing Gaussian spatial noise ($\sigma=0$5050 px) for the I-VT algorithm. Agreement with the noise-free classification drops below 90% at $\sigma=2$ px for Random Pixel, while Waldo maintains agreement above 90% until $\sigma=5$ px. At high noise levels ($\sigma\geq 30$ px), both tasks approach chance-level agreement ($\sim$5–7%), with nearly all samples misclassified as saccades.
Tables S4-S6 compare accuracy and F1 scores between fixed thresholds (optimized at $\sigma=0$) and adaptive thresholds (re-optimized at each noise level) for all three algorithms across both tasks. I-DT with adaptive thresholds demonstrates the greatest robustness, maintaining accuracy above 0.81 even at $\sigma=50$ px, whereas I-AVT shows the steepest degradation, falling to 0.31 accuracy at extreme noise levels.
Figure S5 visualizes F1 scores for fixation and saccade detection across noise levels. Adaptive thresholds (dashed lines) preserve fixation F1 scores substantially better than fixed thresholds (solid lines), with I-DT maintaining fixation F1 >> 0.90 across all noise levels. However, saccade F1 scores decline for all algorithms under both threshold conditions, with adaptive I-DT showing particularly low saccade F1 (<< 0.11) at high noise. Figure S6 reveals the precision-recall trade-off underlying adaptive threshold optimization. Precision for fixation detection remains high across noise levels under adaptive thresholding, but recall decreases substantially at $\sigma>10$ px. This pattern indicates that adaptive thresholds achieve overall accuracy by conservatively classifying uncertain samples as fixations.
The results demonstrate several key patterns regarding noise robustness. Algorithm-specific differences in noise resilience are evident: I-DT with adaptive thresholds maintain the highest accuracy even at extreme noise conditions, while I-AVT shows the steepest degradation, suggesting that angular velocity features are particularly susceptible to noise-induced perturbations. A clear precision-recall trade-off emerges with adaptive thresholds, while adaptive optimization preserves fixation F1 scores, This comes at the cost of saccade detection performance. This strategy may be acceptable in applications prioritizing fixation identification, but problematic when accurate saccade detection is required.